OpenStudy (anonymous):

Factor completely 81x^2 - 4. a. (2x - 9)(2x - 9) b. (2x - 9)(2x + 9) c. (9x - 2)(9x - 2) d. (9x - 2)(9x + 2)

OpenStudy (jdoe0001):

\(\bf 81x^2 - 4\qquad \qquad 81=9^2\qquad \qquad 4=2^2\qquad thus\\ \quad \\ 81x^2 - 4\implies 9^2x^2-2^2\implies (9x)^2-2^2\\ \quad \\ \textit{recall that }\quad a^2-b^2 = (a-b)(a+b) \)

OpenStudy (anonymous):

\[81x ^{2}-4\] Since they are both perfect squares, just take the square root of each. (9x+2)(9x-2) so D

OpenStudy (anonymous):

thank you!cx

OpenStudy (jdoe0001):

yw